OpenStudy (anonymous):

I'm trying to teach myself Algebra over the summer. Can someone explain this problem to me? "The factored expression (m+7) (m-11) is equivalent to the expanded form m^2 - 4m - 77. When m=6, the value of m^2 - 4m - 77 is a multiple of 13. Why is this?

OpenStudy (anonymous):

which part is not clear?

OpenStudy (anonymous):

I don't know why it is always a multiple of 13. That's what I don't understand.

OpenStudy (campbell_st):

just replace m with 6 6^2 - 4*6 - 77 =..... I think you'll find the answer is a multiply of 13..... the other factor is negative

OpenStudy (anonymous):

(m+6) if m=7 is 7+6=13 so we have 13*something So its a multiple of 13

OpenStudy (anonymous):

I still don't get it, I'm sorry.

OpenStudy (anonymous):

the factorized for is (m+6)*(m-11) now if m=7 that is (7+6)*(7-11)=13*(-4)

OpenStudy (anonymous):

that means that it is a multiple of 13 as you can pull that number out

OpenStudy (anonymous):

oh...I get it now, Thank you so much!